OpenStudy (anonymous):

What is the next number in the sequence 1,5, 14,30, 55, 91

OpenStudy (nenadmatematika):

140

OpenStudy (anonymous):

explain

OpenStudy (ash2326):

we have 1, 5, 14, 30, 55, 91 5-1=4=2^2 14-5=9=3^2 30-14=16=4^4 55-30=25=5^2 91-55=36=6^2 so the difference between two consecutive no. is the square of nth term 91 is the 6th term, 7th term is =91+7^2=91+49=140

OpenStudy (anonymous):

hey the 30-14 one doesnt fit in

OpenStudy (ash2326):

30-14=16=4^2 30 is the 4th term, generated by 4^2+14

OpenStudy (ash2326):

sorry I wrote 4^4, it's 4^2

OpenStudy (anonymous):

so ash hw to approach such ques?

OpenStudy (ash2326):

tanu, if you are given a sequence, first check , the difference between consecutive terms, Is it a sequence or is it similar? or check the difference between the differences itself, sometimes it may be a pattern. in some cases alternate terms have a pattern like 1, 2, 5, 6, 9, 10 here 1, 5, 9, generated by adding 4 to the term before the previous 2,6 ,10 are a series

OpenStudy (anonymous):

how do u check for a a.p /g.p

OpenStudy (ash2326):

AP will have a common term added to each term to generate the next term 1, 3, 5, 7 1+2=3 3+2=5 5+2=7 and so on GP will have a common term multiplied to generate the next term 2, 6, 18, 54, 162 6=2*3 18=6*3 54=18*3 162=54*3 and so on note that term added or multipled can either be a positive or negative no. . No restriction on this , can be a irrational no. also

OpenStudy (anonymous):

thanks

OpenStudy (ash2326):

glad to help:)